OpenStudy (anonymous):

A test of the prototype of a new automobile shows that the minimum distance for a controlled stop from 100 km/h to zero is 41 m. Find the acceleration, assuming it to be constant as a fraction of the free-fall acceleration. The acceleration due to gravity is 9.81 m/s 2 . Answer in units of g

OpenStudy (anonymous):

use the equation(s) : d = v x t 41 = 100 x t t = 0.41 of a (minute) use this equation: Acceleration = change in velocity / change in time a = ( final velocity - initial velocity ) / ( change in time) a = ( 0 - 100 ) / (0.41 x 60) ** multiplied by 60 to get time in seconds a = -4.1 ( 1dp) ms^(-2) Is this answer right? its been a while since I touched physics :)